OpenStudy (anonymous):

I need Chemistry help (Connections Unit 6 Lesson 3) 1. A sample of gold (Au) has a mass of 35.12 g. a. Calculate the number of moles of gold (Au) in the sample and record in Table 1. Show your work. b. Calculate the number of atoms of gold (Au) in the sample and record in Table 1. Show your work. 2. A sample of table sugar (sucrose, C12H22O11) has a mass of 1.202 g. a. Calculate the number of moles of C12H22O11 contained in the sample and record in Table 1. Show your work.

OpenStudy (anonymous):

b. Calculate the moles of each element in C12H22O11 and record in Table 1. Show your work. c. Calculate the number of atoms of each type in C12H22O11 and record in Table 1. Show your work. Also, can you help me figure out how to work it out rather than just telling me the answer? I have no idea how to do these.

OpenStudy (anonymous):

1 mole of Au= 196.97g \[1a)35.12g \times \frac{ 1 mole }{ 196.97g }= 0.1783 moles\] \[1b)(6.022\times10^{23})\times0.1783 moles= 1.074\times10^{23} atoms\] \[1 mole C _{12}= 144.132g \] \[1 mole H _{22}= 22.1738g\] \[1 mole O _{11}=175.989g\] \[1 mole C _{12}H _{22}O _{11}= 342.295g\] \[2a)1.202g \times \frac{ 1 mole }{ 342.295g }= 0.0035 moles\] \[2b).0035moles \times12=0.042 moles C\] \[b).0035 moles \times 22= 0.077 moles H\] \[b).0035moles \times11=0.386 moles O\]

OpenStudy (anonymous):

@gusmariewhata Thank you sooo much. Can you help me find the number of atoms in C12H22O11 too? I tried figuring out how you did it with Au, but I honestly have no idea /:

OpenStudy (anonymous):

@gusmariewhata Never mind, I got that part. Thank you so much, again

OpenStudy (anonymous):

@gusmariewhata could you help me figure out how to put this information in the table? i'm confused on where to put everything

OpenStudy (anonymous):

@imuseless

OpenStudy (anonymous):

brittbratt did you end up getting it?

OpenStudy (anonymous):

@sammyjpod could you help?

OpenStudy (anonymous):

Yeah I need help with this too.

OpenStudy (anonymous):

same

OpenStudy (awkwardpanda):

HELP

OpenStudy (anonymous):

Does anyone know how to do the atom part of this????

OpenStudy (anonymous):

hey yall lmao help

OpenStudy (toloaf):

@dalecoopers did you get 2c? I need help. Thanks